# Analyzing News Engagement on Facebook: Tracking Ideological Segregation and News Quality in the Facebook URL Dataset

### v2

@@ 4 Longitudinal Analysis
To quantify these shifts, we identify significant change points by applying segmented linear regression Muggeo (2003); Jekel and Venter (2019) to each individual engagement timeline. Segmented linear regression models the data as a piecewise linear trend, aligning with the observed data and allowing for changes at distinct breakpoints identified as change points.
For most individual engagement metrics, the identified number of change points is two. These are marked by shaded regions in Figure 4 (see the Supplementary Information 5 for more details). We complement these regions with dashed lines indicating dates when actual changes to the Facebook ranking algorithm were reported Mosseri (2018); Hagey and Horwitz (2021); Narayanan (2023). While the change points coincide with algorithmic changes, the extent to which these events are driven by the ranking algorithm or other exogenous factors cannot be determined from this analysis. We discuss this in more detail in Section 5.2.
@@
We conclude this initial analysis by noting that these trends correspond to the raw engagement time series. For example, engagement time series could be normalized by views to reduce common effects across all engagement types. The results of this normalization are shown in the Supplementary Information 6A, Fig. S6, where we observe that the share of active engagement steadily increases over time compared to passive engagement.
@@ 4.2 Measuring Segregation and Low-Quality News Consumption Over Time
For clarity, we merge moderate and strong user political affinities (PPA) into a combined political category $\mathbb{P}$ resulting in four distinct combined categories: conservatives $\mathcal{C}$ (buckets $+2$ and $+1$), liberals $\mathcal{L}$ (buckets $-2$ and $-1$), centrists $\mathcal{N}$ (including only bucket 00), and undefined $\mathcal{D}$ (when the PPA is not defined). Merging strong and moderate buckets does not affect the results obtained.
Formally, let $u$ denote a URL present in the data (we omit the month index to simplify notation) and let $i(u)\in[0,1]$ and $q(u)\in[0,1]$ denote the domain-level ideology and domain-level quality scores of $u$, respectively  [^3] To prevent numerical problems when calculating confidence intervals for the denominators, we normalize the domain ideology scores from 00 to $1$, where 00 indicates the most liberal-leaning domains and $1$ represents the most conservative-leaning domains in our dataset..
@@
Given that we are again working with ratios involving noisy denominators, it is important to carefully evaluate the signal-to-noise ratio (SNR) of the denominator and to determine the uncertainty intervals for each component of the ideology gap. We obtain SNR values exceeding $1\cdot 10^{5}$ and uncertainty intervals for the averages of both $\mathcal{C}$ and $\mathcal{L}$ on the order of $1\cdot 10^{-3}$ (see the Supplementary Information 2B). We therefore can conclude that the metrics are statistically significant. The weighted standard deviations are substantial and vary over time, so we plot them separately when needed.
@@
We also quantify the amount of low-quality content, indicative of misinformation, by calculating the proportion of URLs with a content quality score lower than a threshold $T_{\text{low}}$ Lin et al. (2023):
@@
Figure 5: Evolution of engagement (clicks) in terms of content ideology and content quality. (A) and (D) show the evolution of the weighted averages of the domain ideology (normalized between 0 and 1) and domain quality, respectively, as in Eq (6). The averages are done for each user class (i.e. Conservative, Liberal, Centrist, or without a defined PPA) using clicks as weights. (B) and (E) show the weighted standard deviations related to the ideology and quality averages above, as in Eq (11). (C) represents the ideological gap between the conservative and liberal averages and can serve as a proxy for ideological segregation of news consumption, Eq (12). (F) shows the proportion of clicks directed towards low-quality domains, Eq (13) with $T_{\text{low}}=0.6$. Noise uncertainty intervals are too small to be visually detected in any of the subfigures. For all shown metrics that involve noisy denominators, we find $SNR>10^{5}$. Dashed lines indicate relevant algorithmic updates. Shaded areas are calculated as in the previous Figure 4.
4.3 R5a. Findings Regarding the Ideology Gap
@@
Figure 5C shows the ideology gap between conservative and liberal groups (black line). We observe a gradual decline until the second change point in 2020, where the gap increases sharply. This rise can be attributed to a shift in the average consumption and engagement of conservative-leaning users toward more extreme-leaning domains, as illustrated in Figure 5A. In this period we observe a greater supply of more conservative content as well as an increased engagement with more extreme content from conservative users.
@@
Figure 6: Distribution of Clicks by Domain Quality Over Time. Each column represents the monthly distribution of clicks across domain quality bins (bin width $=0.05$), normalized per month. The quality score bin of $0.58$ serves as a threshold that differentiates high- and medium-quality domains from low-quality ones. For low-quality domains, we highlight three events (A), (B), and (C), which are associated with algorithmic change points. Relevant dates for algorithmic changes are marked with a black line marker. Time intervals marked with a red marker indicate change point regions obtained through piece-wise linear regression fitting.
@@
In Figure 5F we show the change in the proportion of engagement directed at low-quality domains (quality bin lower than the bin centered at $0.58$ in Figure 1 and Figure 6), which fluctuates around $20\%$ and $30\%$. We observe two sudden decreases around change point regions, in the first case after the region (2018-03) and in the second case before the region (2020-03). We also see an immediate increase after the drops in both cases, even though it is steeper in the second region, achieving values similar to the start of the plot, with ${}30\%$ of the engagement directed towards low-quality sources. These two “bounce-back” effects are driven mainly by conservative-learning users. Note that the second bounce-back effect coincides with the Covid-19 pandemic and might be related to its respective information crisis.
In Figure 6 we show the distribution of clicks per month in each domain quality bin. These bins are the same as the ones defined in the aggregate analysis (Figure 3C). We find two clear sections below and above the quality score bin centered at $0.58$, which agrees with the tentative threshold of $\sim 0.6$ for high-quality news given by Lin et al. (2023) and thus is used as a threshold in Figure 5F. The set of domains above this bin, which corresponds to medium and high-quality domains, dominates engagement (most of the clicks fall into that area of the figure), but there still are some low-quality domains that receive comparable engagement.
Regarding the trends in low-quality domains, we highlight three key events in Figure 6 that help us better understand Figure 5F. The first event (A) coincides with the first algorithmic change and is characterized by a sudden drop in engagement with low-quality domains, primarily in domains with quality scores $\approx 0.28$. The second and third events occur during the period of the second algorithmic change: event (B) marks another sharp decline affecting all low-quality source domains, followed by event (C), a rapid rebound driven primarily by domains with quality scores $\approx 0.38$.
@@ 5.2 Contextualization of our results with the Facebook News Feed algorithm
In terms of quality, we believe that a possible explanation for this phenomenon, based on the aggregate analysis, comes from the fact that more extreme users consume more extreme content and, particularly for conservative users, also of lower quality. This, combined with the U-shaped patterns of engagement, suggests that more extreme users are more likely to engage with comments, likes or angers than moderate ones. Hence an algorithmic update that boosts the weights on, say, comments (that are strongly U-shaped), makes posts that are likely to generate comments more visible to users. But because more extreme users are more likely to comment and such users tend to consume more extreme and (in part) also lower quality content, the algorithmic boost will tend to make more extreme and lower quality content more visible. The latter will in turn be more likely to be commented on, making it even more visible, and so on. Related to these questions, Germano et al. (2022) study such feedback loops occurring in response to algorithmic changes, and Chavalarias et al. (2024) also study associated network effects. Going through the feedback loop between algorithm and user engagement, applying the above logic to an update such as the MSI (2018), may well trigger increased engagement but also more polarization (because of the more oppositely extreme content being consumed on both sides) and more low-quality information or misinformation (because of the lower quality content being consumed, especially on the conservative side).
By contrast, the 2020 update seems more difficult to reconcile with the observed patterns, since, unlike the 2018 update, it did not boost the weights on engagement but rather reduced them, especially for the reshares Hagey and Horwitz (2021); Narayanan (2023). We observe reduced engagement but also an increased ideology gap and a stronger prevalence of low-quality sources. One point that may be worth emphasizing is that reshares have a visibly less U-shaped pattern than comments, whose weight remained high in the News Feed ranking algorithm. It is therefore not inconceivable that a reduction in the weight of reshares (not really U-shaped) while maintaining the weights on comments high (strongly U-shaped) may make comments relatively more important and may therefore boost those undesired effects in terms of consumption of more extreme and low-quality sources. Clearly, the precise role of the differences in weights for different engagement patterns and their relation to engagement as well as segregation of news consumption and diffusion of low-quality news and/or misinformation is something that requires access to more detailed information and seems worth exploring further.
@@ 6 Limitations
We conclude by addressing the limitations of this study. Potentially related dataset issues and their implications are described in the Supplementary Information 1A-D.
@@
This study focuses on interactions with some of the most engaged news-related URLs on Facebook from 2017 to 2021, specifically targeting content from the top 1% of the most viewed news domains in the US during this period Buntain et al. (2023). While it is possible to compare part of the analysis with other content from less popular and news-focused sources, it is not possible to study the implications of the ideological or quality bias of those other sources in their consumption. Therefore, this approach provides critical insights into prominent trends in news consumption but may exclude niche content with limited reach. Consequently, the findings should not be generalized to less engaged, less circulated, or non-news content.
@@
Another limitation of this study is that we use ideology and quality scores that are defined at the domain level rather than at the level of specific news URLs. This is well-documented, as using the average partisanship of source audiences provides a relative measure of partisanship Green et al. (2023); González-Bailón et al. (2023); Buntain et al. (2023), rather than an absolute one. In our study, this distinction does not have significant implications for most domains. However, it could become relevant in domains where individual scores exhibit a multimodal distribution. This could result in slightly underestimated measures of polarization and misinformation, as well as slightly overestimated corresponding standard deviations.

### v3

@@ 4 Longitudinal Analysis
To quantify these shifts, we identify significant change points by applying segmented linear regression Muggeo (2003); Jekel and Venter (2019) to each individual engagement timeline. Segmented linear regression models the data as a piecewise linear trend, aligning with the observed data and allowing for changes at distinct change points identified as change points. For most individual engagement metrics, the identified number of change points is two.
Because our temporal resolution is monthly, and because different engagement metrics may respond to change at slightly different rates, we expect some small variation in the precise change point locations across engagement metrics. Instead of treating individual change points in isolation, we apply segmented linear regression independently on multiple time series (representing different engagement metrics) and then aggregate these results to identify broader change regions (i.e., bands of months where multiple metrics indicate potential structural shifts). These regions are marked by shaded regions in Figure 4 and throughout other figures in the manuscript (see the Supplementary Information 4 for more details on the methodology).
We complement these regions with dashed lines indicating the month when actual changes to the Facebook ranking algorithm were reported or expected to have occurred (see Mosseri (2018); Hagey and Horwitz (2021); Narayanan (2023) for the 2018 update), which is associated with precise date; see Narayanan (2023); Dell et al. (2023) for the 2020 update, for which we could not find a precise date but chose March as the most likely month in which it occurred (February and April could not be excluded). While the change points coincide with algorithmic changes, the extent to which these events are driven by the ranking algorithm or other exogenous factors cannot be determined from this analysis. We discuss this in more detail in Section 5.2.
@@
We conclude this initial analysis by noting that these trends correspond to the raw engagement time series. For example, engagement time series could be normalized by views to reduce common effects across all engagement types. The results of this normalization are shown in the Supplementary Information 5A, Fig. S5, where we observe that the share of active engagement steadily increases over time compared to passive engagement.
@@ 4.2 Measuring Segregation and Low-Quality News Consumption Over Time
For clarity, we merge moderate and strong user political affinities (PPA) into a combined political category $\mathbb{P}$ resulting in four distinct combined categories: conservatives $\mathcal{C}$ (buckets +2+2 and +1+1), liberals $\mathcal{L}$ (buckets 2-2 and 1-1), centrists $\mathcal{N}$ (including only bucket 0), and undefined $\mathcal{D}$ (when the PPA is not defined). Merging strong and moderate buckets does not affect the results obtained.
Formally, let $u$ denote a URL present in the data (we omit the month index to simplify notation) and let $i(u)\in[0,1]$ and $q(u)\in[0,1]$ denote the domain-level ideology and domain-level quality scores of $u$, respectively  [^3] To prevent numerical problems when calculating confidence intervals for the denominators, we normalize the domain ideology scores from 0 to $1$, where 0 indicates the most liberal-leaning domains and $1$ represents the most conservative-leaning domains in our dataset..
@@
Given that we are again working with ratios involving noisy denominators, it is important to carefully evaluate the signal-to-noise ratio (SNR) of the denominator and to determine the uncertainty intervals for each component of the ideology gap. We obtain SNR values exceeding $1\cdot 10^{5}$ and uncertainty intervals for the averages of both $\mathcal{C}$ and $\mathcal{L}$ on the order of $1\cdot 10^{-3}$ (see the Supplementary Information 1B). We therefore can conclude that the metrics are statistically significant. The weighted standard deviations are substantial and vary over time, so we plot them separately when needed.
@@
We also quantify the prevalence of content from low-quality sources, which can be indicative of a higher risk of misinformation exposure, by calculating the proportion of URLs with a content quality score lower than a threshold $T_{\text{low}}$ Lin et al. (2023):
@@
4.3 R5a. Findings Regarding Ideological Segregation
@@
Figure 5: Evolution of engagement (clicks) in terms of content ideology and content quality. (A) and (D) show the evolution of the weighted averages of the domain ideology (normalized between 0 and 1) and domain quality, respectively, as in Eq (6). The averages are done for each user class (i.e. Conservative, Liberal, Centrist, or without a defined PPA) using clicks as weights. (B) and (E) show the weighted standard deviations related to the ideology and quality averages above, as in Eq (11). (C) represents the ideological gap between the conservative and liberal averages and can serve as a proxy for ideological segregation of news consumption, Eq (12). (F) shows the proportion of clicks directed towards low-quality domains, Eq (13) with $T_{\text{low}}=0.6$. Noise uncertainty intervals are too small to be visually detected in any of the subfigures. For all shown metrics that involve noisy denominators, we find $SNR>10^{5}$. Moreover, bootstrapped uncertainty intervals for the weighted averages have very small standard deviations (in the order of 0.001), indicating high statistical precision. Dashed lines indicate relevant algorithmic updates. Shaded areas are calculated as in the previous Figure 4.
Figure 6: Distributions of weighted domain ideology and quality scores per user group at key time points. Each violin plot represents the distribution of domain scores (ideology in the top row, quality in the bottom) obtained by applying the engagement-weighted distribution over URLs to the corresponding score function. Horizontal lines indicate the weighted average score for each user group: conservatives (red), liberals (blue), centrists (yellow), and undefined (gray). These averages correspond to the monthly values shown in panels A and D of Figure 5. The figure shows six selected months around the detected change points in engagement.
@@
Figure 5C shows the ideology gap between conservative and liberal groups (black line). We observe a gradual decline, except during the periods associated with the identified change points. In particular, in the second change point in 2020, the gap increases sharply. This rise can be attributed to a shift in the average consumption and engagement of conservative-leaning users toward more extreme-leaning domains, as illustrated in Figure 5A. In this period we observe a greater supply of more conservative content as well as an increased engagement with more extreme content from conservative users.
While our main analyses rely on weighted averages, we acknowledge that alternative summary measures, such as medians or full distributional comparisons, could yield different patterns. However, we verified that using the median instead of the mean results in similar overall trends, particularly regarding the persistent gap between conservative and liberal users and its variations around the algorithmic change points. Moreover, to complement the trends shown in the time series, the top row of Figure 6 presents the full distributions (depicted as violin plots) of the weighted domain ideology scores for each user group at six key time points. Conservative users (in red) consistently exhibit a broader distribution than other groups.
We observe that the initial decrease in the ideology gap is mainly explained by a shift in engagement of conservatives to less extreme right-leaning outlets, creating a noticeable concentration in the mid-right range (compare red distributions between the 1st and 2nd top plot). This decrease stops after the first change point. There is also a decrease in engagement in extreme left-leaning outlets (in blue).
The second change point is characterized by renewed engagement of conservatives with more ideologically extreme conservative domains (compare red distributions between the 5th and 6th top plots). This second shift is not exclusive to conservatives: it is also reflected among center (yellow) and undefined (gray) users, whose engagement distributions show a small but visible shift in mass toward more extreme right-wing outlets. In contrast, liberal users (blue) show a more stable distribution over time, with less structural change across the change points and less engagement with ideologically extreme sources.
Figure 7: Distribution of Clicks Over Time. Relevant dates for algorithmic changes are marked with black line markers. Time intervals marked with a red marker indicate change point regions obtained through piece-wise linear regression. Ticks mark the center of the bin. (top) monthly distribution of clicks across domain ideology bins (bin width =0.05=0.05), normalized per month. (bottom) monthly distribution of clicks across domain quality bins (bin width =0.05=0.05), normalized per month. A quality score threshold at $0.58$ separates high- and medium-quality domains from low-quality ones. For low-quality domains, we highlight three events (A), (B), and (C), which are associated with algorithmic change points.
To close this subsection, the top panel of Figure 7 illustrates the temporal evolution of ideological engagement by showing the monthly distribution of clicks across domain ideology bins. While engagement with non-extreme domains remains relatively stable, we identify two periods of increased engagement with domains close to the ideological center, visible as darker horizontal bands near the middle of the distribution. These periods align with the two local minima of the segregation measure shown in Figure 5C, supporting the trends described earlier in the subsection.
Finally, we remark that the increase in the segregation measure after the second change point, shown in Figure 5C, is clearly reflected in Figure 7 as a sharp rise in clicks to far-right domains. This pattern is consistent with the rise in engagement with ideologically extreme conservative content shown in Figure 6, especially among conservative and, to a lesser extent, undefined and center users.
@@
In Figure 5F we show the change in the proportion of engagement directed at low-quality domains (quality bin lower than the bin centered at $0.58$ in Figure 1), which fluctuates around $20\%$ and $30\%$. We observe two sudden decreases around change point regions, in the first case after the region (2018-03) and in the second case before the region (2020-03). We also see an immediate increase after the drops in both cases, even though it is steeper in the second region, achieving values similar to the start of the plot, with $30\%$ of the engagement directed towards low-quality sources. These two “bounce-back” effects are driven mainly by conservative-learning users. Note that the second bounce-back effect coincides with the Covid-19 pandemic and might be related to its respective information crisis.
Similar to the case of domain ideology, Figure 6 (bottom row) shows violin plots of domain quality distributions for each group for selected months around the main engagement change points. While all distributions are broad, the liberal (and centrist and undefined) groups exhibit a higher concentration of engagement in higher-quality domains compared to conservatives.
We observe two main changes in the low-quality region of domain quality scores which appear temporally related to the two minima of Figure 5F. The first, linked to the first change point, is characterized by the disappearance of a set of outlets with quality scores around $0.28$. This shift appears mainly in the center/undefined ideology user classes (compare 1st and 2nd bottom violin plots). The second change, associated with the second change point, is marked by a progressive increase in engagement with domains scoring around $0.38$, as shown in the last three panels of the bottom row.
In Figure 7 (bottom) we show the distribution of clicks per month in each domain quality bin. The two distinct regions below and above the quality score bin centered at $0.58$ support our choice of a threshold around $0.6$ for trustworthy news, as recommended by NewsGuard Lin et al. (2023). The set of domains above this bin, which corresponds to medium and high-quality domains, dominates engagement (most of the clicks fall into that area of the figure), but there still are some low-quality domains that receive comparable engagement.
Regarding the trends in low-quality domains, we highlight three key events in Figure 7 that help us better understand Figures 5F and  6 (bottom row). The first event (A) coincides with the first algorithmic change and is characterized by a sudden drop in engagement with domains with quality scores around $0.28$. The second and third events occur during the period of the second algorithmic change: event (B) marks another sharp decline affecting all low-quality source domains, followed by event (C), a rapid rebound driven primarily by domains with quality scores around $0.38$.
@@ 5.2 Contextualization of our results with the Facebook News Feed algorithm
In terms of quality, we believe that a possible explanation for this phenomenon, based on the aggregate analysis, comes from the fact that more extreme users consume more extreme content and, particularly for conservative users, also of lower quality. This, combined with the U-shaped patterns of engagement, suggests that more extreme users are more likely to engage with comments, likes or angers than moderate ones. Hence an algorithmic update that boosts the weights on, say, comments (that are strongly U-shaped), makes posts that are likely to generate comments more visible to users. But because more extreme users are more likely to comment and such users tend to consume more extreme and (in part) also lower quality content, the algorithmic boost will tend to make more extreme and lower quality content more visible. The latter will in turn be more likely to be commented on, making it even more visible, and so on. Related to these questions, Germano et al. (2022) study such feedback loops occurring in response to algorithmic changes, and Chavalarias et al. (2024) also study associated network effects. Going through the feedback loop between algorithm and user engagement, applying the above logic to an update such as the MSI (2018), may well trigger increased engagement but also more ideological segregation in the distribution of that engagement (because of the more oppositely extreme content being consumed on both sides) and more low-quality information (because of the lower quality content being consumed, especially on the conservative side).
By contrast, the 2020 update seems more difficult to reconcile with the observed patterns, since, unlike the 2018 update, it did not boost the weights on engagement but rather reduced them, especially for the reshares Hagey and Horwitz (2021); Narayanan (2023). We observe reduced engagement but also an increased ideology gap and a stronger prevalence of low-quality sources. One point that may be worth emphasizing is that reshares have a visibly less U-shaped pattern than comments, whose weight remained high in the News Feed ranking algorithm. It is therefore not inconceivable that a reduction in the weight of reshares (not really U-shaped) while maintaining the weights on comments high (strongly U-shaped) may make comments relatively more important and may therefore boost those undesired effects in terms of consumption of more extreme and low-quality sources. Clearly, the precise role of the differences in weights for different engagement patterns and their relation to engagement as well as segregation of news consumption and diffusion of low-quality news is something that requires access to more detailed information and seems worth exploring further.
@@ 6 Limitations
We conclude by addressing the limitations of this study in Subsection 6.1. Potentially related dataset issues and their implications are described in Subsection 6.2.
6.1 Study limitations
@@
This work focuses on interactions with some of the most engaged news-related URLs on Facebook from 2017 to 2021, specifically targeting content from the top 1% of the most viewed news domains in the US during this period Buntain et al. (2023), which accounts for over $77\%$ of the total URL views on Facebook in the US from 2017 to 2020. Given the heavy-tailed distribution of engagement on social media platforms, this cutoff is not arbitrary but reflects a pragmatic balance between coverage and signal quality. Including low-engagement domains would introduce substantial noise due to the platform’s differential privacy constraints, as such domains typically hover near the dataset’s inclusion threshold and offer little usable signal. While it is possible to compare part of the analysis with other content from less popular or non-news sources, it is not feasible to meaningfully assess the ideological or quality bias of such content. Therefore, our approach provides critical insights into prominent trends in news consumption but may exclude niche content with limited reach. Consequently, the findings should not be generalized to less engaged, less circulated, or non-news content.
Furthermore, because the data and engagement behaviors analyzed are specific to Facebook, our findings should not be generalized to other platforms or to the broader online information ecosystem without caution. Different social media platforms have distinct affordances, audiences, and content ranking mechanisms, which may shape ideological and quality-related engagement in different ways.
@@
Another limitation of this study is that we use ideology and quality scores that are defined at the domain level rather than at the level of specific news URLs. This is well-documented, as using the average partisanship of source audiences provides a relative measure of partisanship Green et al. (2025); González-Bailón et al. (2023); Buntain et al. (2023), rather than an absolute one. In our study, this distinction does not have significant implications for most domains. However, it could become relevant in domains where individual scores exhibit a multimodal distribution. This could result in slightly underestimated measures of the ideological gap, as well as slightly overestimated corresponding standard deviations.
Similarly, the domain-level quality scores used in this study are composite metrics that aggregate multiple dimensions (e.g. factual accuracy, transparency, and editorial standards) into a single reliability score Lin et al. (2023). While this provides extensive coverage, it limits interpretability, as a low score may reflect different underlying shortcomings across domains. As such, our analysis captures general trends regarding the reputation of domains but cannot disentangle specific quality-related mechanisms at the content level.
A related limitation of our study lies in its reliance on aggregate engagement data, which precludes us from analyzing individual-level media consumption behaviors. This raises the possibility that some of the trends we observe—particularly the growing ideological gap and engagement with low-quality content—may be driven by a relatively small group of highly active users. This concern has been raised by Guess Guess (2021), who found that while a small subset of users consistently engaged with highly partisan and ideologically skewed online media content, most individuals exhibit far more moderate and diverse media diets. Our results, while robust at the aggregate level, must therefore be interpreted with caution, as they may reflect the outsized influence of this hyper-engaged minority. This phenomenon of aggregation bias is an important consideration in the interpretive validity of our findings.
Moreover, our metrics not only do not make claims about individual-level ideological shifts, but also do not address affective polarization, or user attitudes toward opposing groups. Nor do we interpret our findings as direct evidence of increasing societal polarization. Rather, we study how user engagement behavior, which is shaped in part by algorithmic design choices, relates to the ideological distribution of news content consumed on the platform.
@@
6.2 Potential Dataset Issues and their Impact in this Study
The 100 shares cutoff— The dataset only includes URLs that were shared publicly at least $100$ times a month (plus Laplacian noise) on Facebook Messing et al. (2020). While this censoring can generate bias Allen et al. (2021), it should not particularly affect our study, since we are already applying a filtering on news from popular domains (see Subsection  2.1). However, this bias may skew towards viral posts within those domains, as well as a particular subset of users who engage with such content. Also, since we do not have access to individual-level data, we may be missing the effects of small population segments that may contribute to spreading low-quality information through reshares.
Impact of the Privacy Protection— The Facebook URL dataset is protected with differential privacy Messing et al. (2020); Evans and King (2023) by perturbing URL counts with zero-centered Gaussian noise with variance depending on the interaction type, to prevent tracing actions to specific users or URLs.
Evans and King Evans and King (2023) show that ignoring such differential privacy protections can lead to unpredictable biased results. They provide corrections in feasible scenarios, especially for linear models. For example, the computation of uncertainty intervals in engagement counts is detailed in the Supplementary Information 1A. For more complex models that involve, e.g., weighted average metrics, there are no available corrections in closed form. For situations involving a noisy denominator, such as the ones in this work, Evans and King Evans and King (2021) and Buntain et al. Buntain et al. (2023) show that for high enough signal-to-noise ratio estimates of the denominator, it is possible to ensure reliable metrics.
[…]
Past issues with the Dataset— The dataset has encountered occasional data integrity issues 444 See for example: https://www.nytimes.com/live/2020/2020-election-misinformation-distortions#facebook-sent-flawed-data-to-misinformation-researchers.. A significant discrepancy was identified on September 10, 2021, where data from US-based users with undefined PPA, which account for more than half of the engagement in the platform, was missing. This issue was detected after researchers had already begun using the dataset and was reportedly retroactively corrected Allen et al. (2022). Such lapses highlight the challenges of relying on datasets managed by platforms with limited transparency in their data collection processes.